# A physics-compliant diagonal representation for wireless channels parametrized by beyond-diagonal reconfigurable intelligent surfaces

### v1

@@ II-B Conventional perspective
II-C Alternative perspective
Alternatively, we can begin by evaluating the cascade of RE and LC, whose auxiliary ports associated with the tunable lumped elements will then be terminated by IL. The cascade of RE and LC is characterized by
$\mathbf{S}^{\mathrm{casc}}=\begin{bmatrix}\mathbf{S}^{\mathrm{casc}}_{\mathcal{AA}}&\mathbf{S}^{\mathrm{casc}}_{\mathcal{AC}}\\ \mathbf{S}^{\mathrm{casc}}_{\mathcal{CA}}&\mathbf{S}^{\mathrm{casc}}_{\mathcal{CC}}\end{bmatrix}\in\mathbb{C}^{N_{\mathrm{casc}}\times N_{\mathrm{casc}}},$ (7)
@@ II-C Alternative perspective
where $N_{\mathrm{casc}}=N_{\mathrm{A}}+N_{\mathrm{C}}$, and $\mathcal{A}$ and $\mathcal{C}$ denote the sets of port indices associated with antennas and tunable lumped elements, respectively. The analytical expression for $\mathbf{S}^{\mathrm{casc}}$ is obtained via the well-established “Redheffer star product” [41, 42, 43, 44]:
$\begin{split}&{\mathbf{S}}^{\mathrm{casc}}_{\mathcal{A}\mathcal{A}}=\mathbf{S}^{\mathrm{RE}}_{\mathcal{A}\mathcal{A}}-\mathbf{S}^{\mathrm{RE}}_{\mathcal{A}\mathcal{S}}\mathbf{S}^{\mathrm{LC}}_{\bar{\mathcal{S}}\bar{\mathcal{S}}}\mathbf{X}_{1}\mathbf{S}^{\mathrm{RE}}_{\mathcal{S}\mathcal{A}},\\ &{\mathbf{S}}^{\mathrm{casc}}_{\mathcal{A}\mathcal{C}}=-\mathbf{S}^{\mathrm{RE}}_{\mathcal{A}\mathcal{S}}\mathbf{X}_{2}\mathbf{S}^{\mathrm{LC}}_{\bar{\mathcal{S}}\mathcal{C}},\\ &{\mathbf{S}}^{\mathrm{casc}}_{\mathcal{C}\mathcal{A}}=-\mathbf{S}^{\mathrm{LC}}_{\mathcal{C}\bar{\mathcal{S}}}\mathbf{X}_{1}\mathbf{S}^{\mathrm{RE}}_{\mathcal{S}\mathcal{A}},\\ &{\mathbf{S}}^{\mathrm{casc}}_{\mathcal{C}\mathcal{C}}=\mathbf{S}^{\mathrm{LC}}_{\mathcal{C}\mathcal{C}}-\mathbf{S}^{\mathrm{LC}}_{\mathcal{C}\bar{\mathcal{S}}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{S}\mathcal{S}}\mathbf{X}_{2}\mathbf{S}^{\mathrm{LC}}_{\bar{\mathcal{S}}\mathcal{C}},\end{split}$ (8)
@@
This expression for the cascade of RE and LC is mathematically more complex than that for the cascade of LC and IL provided in the previous section because both RE and LC have “free” (unconnected) ports after the connection of RE and LC whereas only LC but not IL has free ports after the connection of LC and IL [44].
Given $\mathbf{S}^{\mathrm{casc}}$, we can then determine $\tilde{\mathbf{S}}$ as follows:
$\tilde{\mathbf{S}}=\mathbf{S}^{\mathrm{casc}}_{\mathcal{AA}}+\mathbf{S}^{\mathrm{casc}}_{\mathcal{AC}}\left(\left(\mathbf{S}^{\mathrm{IL}}\right)^{-1}-\mathbf{S}^{\mathrm{casc}}_{\mathcal{CC}}\right)^{-1}\mathbf{S}^{\mathrm{casc}}_{\mathcal{CA}}.$ (10)
The wireless channel matrix is then again obtained via Eq. (3). The development outlined in this subsection hence constitutes the physics-compliant diagonal representation for wireless channels parametrized by BD-RIS because $\mathbf{S}^{\mathrm{IL}}$ in Eq. (10) is always diagonal.
II-D Comparison
@@
Obviously, the expression for $\tilde{\mathbf{S}}$ from Eq. (10) must be equal to the expression for $\tilde{\mathbf{S}}$ from Eq. (2) since we did not make any assumptions to obtain either of these equations. It is instructive to compare the role of corresponding quantities in Eq. (2) and Eq. (10):
$\mathbf{S}^{\mathrm{RE}}\leftrightarrow\mathbf{S}^{\mathrm{casc}},$ (11a)
$\mathbf{S}^{\mathrm{L}}\leftrightarrow\mathbf{S}^{\mathrm{IL}},$ (11b)
$\mathcal{S}\leftrightarrow\mathcal{C}.$ (11c)
@@
In the conventional perspective, the primary ports are those comprised in $\mathcal{A}\cup\mathcal{S}$, and the ports of RE from $\mathcal{S}$ are terminated by $\mathbf{S}^{\mathrm{L}}$ which is generally “beyond-diagonal”. In the alternative perspective, the primary ports are those comprised in $\mathcal{A}\cup\mathcal{C}$, and the ports of the cascade of RE and LC from $\mathcal{C}$ are terminated by $\mathbf{S}^{\mathrm{IL}}$ which is always diagonal.
III Determining $\mathbf{S}^{\mathrm{LC}}$
@@
At the core of our theoretical developments in the previous section is $\mathbf{S}^{\mathrm{LC}}$, the scattering matrix characterizing the static load circuit which connects the RIS elements to the tunable individual loads. The existence of $\mathbf{S}^{\mathrm{LC}}$ has to date not been brought up explicitly in the BD-RIS literature, to the best of our knowledge, such that we dedicate this section to discussing how to determine the values of $\mathbf{S}^{\mathrm{LC}}$ analytically for canonical examples of ideal load circuits, as well as how it could be determined in numerical full-wave simulations and to what extent it could be determined experimentally. At the end of this section, we go on to argue that, in fact, operationally there is no need to explicitly know $\mathbf{S}^{\mathrm{LC}}$ and that what we need to know or estimate is $\mathbf{S}^{\mathrm{casc}}$ but not its breakdown into $\mathbf{S}^{\mathrm{RE}}$ and $\mathbf{S}^{\mathrm{LC}}$.
@@ III-A Theoretical canonical examples
Theoretically studied tunable circuits that terminate RIS elements in BD-RIS typically include tunable lumped elements that are connected without loss, delay and dispersion to each other and/or the RIS element auxiliary ports and/or ground. Under these simplifying assumptions, the multi-port network description of the circuit can be determined analytically.
Thus far, existing theoretical studies of BD-RIS directly obtained $\mathbf{S}^{\mathrm{L}}$ without explicitly identifying $\mathbf{S}^{\mathrm{LC}}$ and $\mathbf{S}^{\mathrm{IL}}$. More specifically, as detailed in Eq. (13) of Ref. [7] (based on Eq. (6) of Ref. [45]), the admittance representation $\mathbf{Y}^{\mathrm{L}}$ of the circuit is obtained in closed form, which can then be transformed to $\mathbf{S}^{\mathrm{L}}$, if desired.
@@
Figure 2: (A) Ideal 2-port T network. (B) The three impedances in (A) are replaced by auxiliary ports, yielding a 5-port network whose scattering matrix can be determined analytically, see the Appendix and Eq. (12). (C) Ideal 2-port $\pi$ network. (D) The three impedances in (C) are replaced by auxiliary ports, yielding a 5-port network whose scattering matrix can be determined analytically, see Eq. (13).
Using basic circuit analysis, as detailed in the Appendix, we readily obtain the $5\times 5$ scattering matrix of the circuit from Fig. 2B:
$\mathbf{S}^{\mathrm{LC,T}}=\begin{bmatrix}0.25&0.25&-0.75&-0.25&0.5\\ 0.25&0.25&0.25&0.75&0.5\\ -0.75&0.25&0.25&-0.25&0.5\\ -0.25&0.75&-0.25&0.25&-0.5\\ 0.5&0.5&0.5&-0.5&0\end{bmatrix}.$ (12)
We have verified this result with three sanity checks, as detailed in the Appendix.
@@
$\mathbf{S}^{\mathrm{LC,\pi}}=\begin{bmatrix}-0.25&0.25&-0.5&0.75&0.25\\ 0.25&-0.25&0.5&0.25&0.75\\ -0.5&0.5&0&-0.5&0.5\\ 0.75&0.25&-0.5&-0.25&0.25\\ 0.25&0.75&0.5&0.25&-0.25\end{bmatrix}.$ (13)
@@
The same exercise can be repeated for more complex ideal reconfigurable impedance networks such as the one from Fig. 2(b) in Ref. [7] but manual execution of the required circuit analysis rapidly becomes tiresome; already for the example from Fig. 2(b) in Ref. [7] the dimensions of the sought-after $\mathbf{S}^{\mathrm{LC}}$ would be $14\times 14$. Therefore, for studies of large and/or realistic (as opposed to ideal) reconfigurable impedance networks, the numerical or experimental approaches described in the next two subsections are more suitable.

### v2

@@ II-A Background on Multi-Port Network Models for RIS-Parametrized Channels
The principle of this derivation of the widespread simplified cascaded model from the physics-compliant model was first proposed for D-RIS based on PhysFad in Ref. [20]; here, we have transposed it to multi-port network theory using scattering parameters, and extended it to the BD-RIS case.
II-B Key Insight
[…]
Our key insight is that the load circuit (L) of a BD-RIS, irrespective of its architectural details, can always be separated into a static part (SLC) and a set of tunable individual lumped elements (IL). We denote the number of tunable elements within the BD-RIS load circuit by $N_{\mathrm{C}}$; typically, $N_{\mathrm{C}}>N_{\mathrm{S}}$ but our theory holds equally for $N_{\mathrm{C}}\leq N_{\mathrm{S}}$. It follows that LC must be amenable to a description as a multi-port network with $N_{\mathrm{SLC}}=N_{\mathrm{S}}+N_{\mathrm{C}}$ ports, of which $N_{\mathrm{S}}$ are connected to the RE’s auxiliary ports associated with RIS elements and $N_{\mathrm{C}}$ are terminated by the individual tunable loads comprised in IL. Consequently, L is the result of cascade-loading SLC with IL, analogous to an RE parametrized by a D-RIS. By analogy with Eq. (1), it directly follows that
$\mathbf{S}^{\mathrm{L}}=\mathbf{S}^{\mathrm{SLC}}_{\bar{\mathcal{S}}\bar{\mathcal{S}}}+\mathbf{S}^{\mathrm{SLC}}_{\bar{\mathcal{S}}\mathcal{C}}\left(\left(\mathbf{S}^{\mathrm{IL}}\right)^{-1}-\mathbf{S}^{\mathrm{SLC}}_{\mathcal{CC}}\right)^{-1}\mathbf{S}^{\mathrm{SLC}}_{\mathcal{C}\bar{\mathcal{S}}},$ (5)
where $\mathbf{S}^{\mathrm{SLC}}\in\mathbb{C}^{N_{\mathrm{SLC}}\times N_{\mathrm{SLC}}}$ is the scattering matrix of SLC and $\mathbf{S}^{\mathrm{IL}}\in\mathbb{C}^{N_{\mathrm{C}}\times N_{\mathrm{C}}}$ is the scattering matrix of IL. $\mathbf{S}^{\mathrm{IL}}$ is diagonal; its $i$th entry is the reflection coefficient of the $i$th tunable lumped element within the load circuit. $\bar{\mathcal{S}}$ and $\mathcal{C}$ denote the sets of port indices of SLC associated with RIS elements and tunable individual loads, respectively.
At this stage, the following three remarks are important:
  • The physics-compliant decomposition of the BD-RIS load circuit into static parts and tunable lumped elements, as presented in Eq. (5), underpins the key insight of this paper. This decomposition was thus far overlooked in the BD-RIS literature.
[…]
  • We dedicate Sec. III to explaining how to determine $\mathbf{S}^{\mathrm{SLC}}$ in any given theoretical, numerical or experimental study involving BD-RIS. However, operationally, it is usually not necessary to explicitly determine $\mathbf{S}^{\mathrm{SLC}}$, as seen in our case study in Sec. IV.
Given the insight that L is a cascade of SLC and IL, it follows that the end-to-end BD-RIS-parametrized wireless channel arises in fact from the chain cascade of three multi-port systems: RE, SLC and IL. A detailed illustration thereof is provided in Fig. 1. Conventionally, the separation of L into SLC and IL is not considered. In other words, the conventional representation of BD-RIS-parametrized channels implicitly assumes that SLC and IL have already been cascaded, yielding L, such that one is left with the cascade loading of RE by L. In this representation, one works directly with Eq. (2) in which $\mathbf{S}^{\mathrm{L}}$ is a “beyond-diagonal” matrix.
The alternative representation of BD-RIS-parametrized channels proposed in this paper starts instead by evaluating the cascade of RE and LC, yielding a system (named K) encapsulating all static parts within the chain cascade; then, the auxiliary ports of K associated with the tunable lumped elements are terminated by IL. K is characterized by
$\mathbf{S}^{\mathrm{K}}=\begin{bmatrix}\mathbf{S}^{\mathrm{K}}_{\mathcal{AA}}&\mathbf{S}^{\mathrm{K}}_{\mathcal{AC}}\\ \mathbf{S}^{\mathrm{K}}_{\mathcal{CA}}&\mathbf{S}^{\mathrm{K}}_{\mathcal{CC}}\end{bmatrix}\in\mathbb{C}^{N_{\mathrm{K}}\times N_{\mathrm{K}}},$ (6)
@@ II-B Key Insight
where $N_{\mathrm{K}}=N_{\mathrm{A}}+N_{\mathrm{C}}$. The analytical expression for $\mathbf{S}^{\mathrm{K}}$ is obtained via the well-established “Redheffer star product” [45, 46, 47, 44]:
$\begin{split}&{\mathbf{S}}^{\mathrm{K}}_{\mathcal{A}\mathcal{A}}=\mathbf{S}^{\mathrm{RE}}_{\mathcal{A}\mathcal{A}}-\mathbf{S}^{\mathrm{RE}}_{\mathcal{A}\mathcal{S}}\mathbf{S}^{\mathrm{LC}}_{\bar{\mathcal{S}}\bar{\mathcal{S}}}\mathbf{X}_{1}\mathbf{S}^{\mathrm{RE}}_{\mathcal{S}\mathcal{A}},\\ &{\mathbf{S}}^{\mathrm{K}}_{\mathcal{A}\mathcal{C}}=-\mathbf{S}^{\mathrm{RE}}_{\mathcal{A}\mathcal{S}}\mathbf{X}_{2}\mathbf{S}^{\mathrm{LC}}_{\bar{\mathcal{S}}\mathcal{C}},\\ &{\mathbf{S}}^{\mathrm{K}}_{\mathcal{C}\mathcal{A}}=-\mathbf{S}^{\mathrm{LC}}_{\mathcal{C}\bar{\mathcal{S}}}\mathbf{X}_{1}\mathbf{S}^{\mathrm{RE}}_{\mathcal{S}\mathcal{A}},\\ &{\mathbf{S}}^{\mathrm{K}}_{\mathcal{C}\mathcal{C}}=\mathbf{S}^{\mathrm{LC}}_{\mathcal{C}\mathcal{C}}-\mathbf{S}^{\mathrm{LC}}_{\mathcal{C}\bar{\mathcal{S}}}\mathbf{S}^{\mathrm{RE}}_{\mathcal{S}\mathcal{S}}\mathbf{X}_{2}\mathbf{S}^{\mathrm{LC}}_{\bar{\mathcal{S}}\mathcal{C}},\end{split}$ (7)
@@
This expression for $\mathbf{S}^{\mathrm{K}}$ is mathematically more complex than that for $\mathbf{S}^{\mathrm{L}}$ in Eq. (5) because both RE and LC have “free” (unconnected) ports after their connection whereas only LC but not IL has free ports after their connection [44].
Given $\mathbf{S}^{\mathrm{K}}$ and $\mathbf{S}^{\mathrm{IL}}$, the standard “cascade loading” formula directly yields $\tilde{\mathbf{S}}$; specifically, analogous to Eq. (1), we obtain
$\tilde{\mathbf{S}}=\mathbf{S}^{\mathrm{K}}_{\mathcal{AA}}+\mathbf{S}^{\mathrm{K}}_{\mathcal{AC}}\left(\left(\mathbf{S}^{\mathrm{IL}}\right)^{-1}-\mathbf{S}^{\mathrm{K}}_{\mathcal{CC}}\right)^{-1}\mathbf{S}^{\mathrm{K}}_{\mathcal{CA}}$ (9)
and hence
$\mathbf{H}=\left[\tilde{\mathbf{S}}\right]_{\mathcal{RT}}=\mathbf{S}^{\mathrm{K}}_{\mathcal{RT}}+\mathbf{S}^{\mathrm{K}}_{\mathcal{RC}}\left(\left(\mathbf{S}^{\mathrm{IL}}\right)^{-1}-\mathbf{S}^{\mathrm{K}}_{\mathcal{CC}}\right)^{-1}\mathbf{S}^{\mathrm{K}}_{\mathcal{CT}}.$ (10)
Importantly, $\mathbf{S}^{\mathrm{IL}}$ is a diagonal matrix such that the mathematical structure of Eq. (10) for BD-RIS is analogous to that of Eq. (1) in the D-RIS case. This is our key result.
@@
Obviously, the expressions for $\mathbf{H}$ from Eq. (10) must be equal to that from Eq. (2) since we did not make any assumptions to obtain either of these equations. It is instructive to compare the role of corresponding quantities in these equations:
[…]
$\mathbf{S}^{\mathrm{RE}}\leftrightarrow\mathbf{S}^{\mathrm{K}},$ (11b)
$\mathbf{S}^{\mathrm{L}}\leftrightarrow\mathbf{S}^{\mathrm{IL}},$ (11c)
$\mathcal{S}\leftrightarrow\mathcal{C}.$ (11d)
@@
In the conventional representation, the primary ports are those comprised in $\mathcal{A}\cup\mathcal{S}$, and the ports of RE from $\mathcal{S}$ are terminated by $\mathbf{S}^{\mathrm{L}}$ which is generally “beyond-diagonal”. In our proposed alternative representation, the primary ports are those comprised in $\mathcal{A}\cup\mathcal{C}$, and the ports of K within $\mathcal{C}$ are terminated by $\mathbf{S}^{\mathrm{IL}}$ which is always diagonal. We reiterate that the conventional representation and our alternative representation are both physics-compliant and fully equivalent. However, the alternative representation allows us to directly apply physics-compliant D-RIS algorithms to BD-RIS scenarios, as we demonstrate in Sec. IV.
II-C Transposition to Coupled-Dipole Model (PhysFad) for RIS-Parametrized Channels
Thus far, we have only discussed circuit-theory-based physics-compliant modeling of BD-RIS parametrized channels in terms of scattering parameters. For the D-RIS case, an alternative physics-compliant approach based on a coupled-dipole formalism exists. It was introduced as PhysFad [17] and constituted the basis of the first experimental validation of a physics-compliant channel model for D-RIS-parametrized rich-scattering channels as well as various frugal physics-compliant end-to-end channel estimation protocols [12]. In this subsection, we first review the basics of PhysFad (Sec. II-C1). Then, we establish the operational equivalence between PhysFad and scattering-parameter-based circuit models for D-RIS (Sec. II-C2). Finally, combining Sec. II-C2 with our diagonal scattering-parameter-based representation of BD-RIS-parametrized channels from Sec. II-B, we will describe BD-RIS-parametrized channels within the PhysFad formalism (Sec. II-C3).
II-C1 Background on PhysFad for D-RIS
PhysFad assigns a dipole to each of the $N=N_{\mathrm{A}}+N_{\mathrm{S}}$ ports of the primary wireless entities: $N_{\mathrm{A}}$ antenna ports and $N_{\mathrm{S}}$ auxiliary ports associated with each RIS elements’ tunable lumped elements. The $i$th dipole is characterized by its polarizability $\alpha_{i}$, which is tunable in the case of the tunable lumped elements. In the compact formulation of PhysFad [18, 19, 48], the $i$th and $j$th dipoles are coupled to each other via the background Green’s function $G_{ij}=G_{ji}$ which takes into account all scattering effects within the radio environment (scattering objects of any size and material, including structural scattering of antennas and RIS elements).
Then, an interaction matrix $\mathbf{W}\in\mathbb{C}^{N\times N}$ is set up, composed of the sum of a diagonal matrix containing the inverse polarizabilities and a symmetric matrix containing the background Green’s functions:
$W_{i,j}=\begin{cases}\alpha_{i}^{-1},&i=j\\ -G_{ij},&i\neq j\end{cases}.$ (12)
In PhysFad, we define the RIS configuration as the difference between some choice of reference inverse polarizability $\alpha_{0}^{-1}$ and the RIS elements’ inverse polarizabilities:
$\mathbf{c}=\left[\alpha_{0}^{-1}-\alpha_{i}^{-1}\ |\ i\in\mathcal{S}\right].$ (13)
A convenient partioned representation of $\mathbf{W}(\mathbf{c})$ follows:
[…]
The end-to-end channel matrix is finally identified as the $\mathcal{RT}$ block of the inverse of this interaction matrix [17]:
$\mathbf{H}_{\mathrm{PF}}(\mathbf{c})=\left[\mathbf{W}^{-1}(\mathbf{c})\right]_{\mathcal{RT}}.$ (15)
Whereas circuit theory treats systems as black boxes, PhysFad’s coupled-dipole approach describes the physical wave-matter interactions inside the systems, enabling insights that are not accessible via circuit theory such as the method for performing conjugate beamforming without pilot exchange purely based on user motion that was proposed and experimentally demonstrated in Ref. [48].
II-C2 Operational Equivalence
We now elaborate on the operational equivalence between the scattering-parameter-based circuit representation used in Sec. II-A and the PhysFad representation described in Sec. II-C1. By “operational equivalence”, we mean that PhysFad has the same mathematical structure as the multi-port network formulation, such that end-to-end channel estimation and optimization for D-RIS are equivalent with both formulations.
We define $\mathbf{\Omega}^{0}=\mathbf{W}^{-1}(\mathbf{c}=\mathbf{0})$ and, after a few basic linear algebra steps detailed in Appendix B, we obtain
$\mathbf{H}_{\mathrm{PF}}(\mathbf{c})=\mathbf{\Omega}^{0}_{\mathcal{RT}}+\mathbf{\Omega}^{0}_{\mathcal{RS}}\left(\mathbf{\Phi}-\mathbf{\Omega}^{0}_{\mathcal{SS}}\right)^{-1}\mathbf{\Omega}^{0}_{\mathcal{ST}}.$ (16)
The mathematical structure of Eq. (16) based on PhysFad is identical to that of Eq. (2) based on multi-port network theory, and the following correspondences are apparent:
[…]
$\mathbf{S}^{\mathrm{RE}}\leftrightarrow\mathbf{\Omega}^{0},$ (17b)
$\mathbf{S}^{\mathrm{L}}\leftrightarrow\mathbf{\Phi}.$ (17c)
II-C3 PhysFad for BD-RIS
By comparing Eq. (11d) and Eq. (17c), it is immediately clear that our proposed diagonal representation for BD-RIS-parametrized channels from Sec. II-B allows us also to model BD-RIS within PhysFad. The physical interpretation is as follows: The dipoles in the compact formulation of PhysFad are associated with the ports of the primary wireless entities. In a BD-RIS-parametrized channel, these are the $N_{\mathrm{A}}$ antenna ports and the $N_{\mathrm{C}}$ auxiliary ports associated with the tunable individual lumped elements within the load circuit. Hence, the representation of a BD-RIS-parametrized radio environment in PhysFad involves $N_{\mathrm{K}}$ dipoles, of which the $N_{\mathrm{C}}$ dipoles associated with tunable lumped elements in the load circuit have tunable polarizabilities. The background Green’s functions between these $N_{\mathrm{K}}$ dipoles account for the potentially highly complex wave scattering within K (cascade of RE and SLC).
The rest of this paper uses our diagonal representation in terms of scattering parameters within the circuit-theoretic model described in Sec. II-B. However, because of the direct mathematically equivalent structure evidenced in Eq. (17c), it is obvious that all details of our experimentally grounded case study on end-to-end channel estimation and optimization in Sec. IV would be exactly identical if the PhysFad formalism from this subsection was used instead.
III Determining $\mathbf{S}^{\mathrm{SLC}}$
@@
At the core of our theoretical developments in Sec. II-B is $\mathbf{S}^{\mathrm{SLC}}$, the scattering matrix characterizing the static parts of the load circuit which connect the RIS elements to the tunable individual loads. The notion of $\mathbf{S}^{\mathrm{SLC}}$ has to date not been brought up explicitly in the BD-RIS literature, to the best of our knowledge, such that we dedicate this section to discussing how to determine $\mathbf{S}^{\mathrm{SLC}}$ in various scenarios that can arise in theoretical, numerical or experimental studies of BD-RIS. At the end of this section, we go on to argue that, in fact, operationally there is no need to explicitly know $\mathbf{S}^{\mathrm{SLC}}$; only an estimate of $\mathbf{S}^{\mathrm{K}}$ is needed but not its breakdown into $\mathbf{S}^{\mathrm{RE}}$ and $\mathbf{S}^{\mathrm{SLC}}$, as evidenced in our case study in Sec. IV.
@@ III-A Theoretical canonical examples
Theoretically studied tunable circuits that terminate RIS elements in BD-RIS typically include tunable lumped elements that are connected without loss, delay and dispersion to each other and/or the RIS element auxiliary ports and/or ground. Under these simplifying assumptions, $\mathbf{S}^{\mathrm{SLC}}$ can be determined analytically.
Thus far, existing theoretical studies of BD-RIS directly obtained $\mathbf{S}^{\mathrm{L}}$ without explicitly identifying $\mathbf{S}^{\mathrm{SLC}}$ and $\mathbf{S}^{\mathrm{IL}}$. More specifically, as detailed in Eq. (13) of Ref. [7] (based on Eq. (6) of Ref. [49]), the admittance representation $\mathbf{Y}^{\mathrm{L}}$ of the circuit is obtained in closed form, which can then be transformed to $\mathbf{S}^{\mathrm{L}}$, if desired.
@@
Figure 2: (A) Ideal 2-port T network. (B) The three impedances in (A) are replaced by auxiliary ports, yielding a 5-port network whose scattering matrix can be determined analytically, see Appendix A and Eq. (18). (C) Ideal 2-port $\pi$ network. (D) The three impedances in (C) are replaced by auxiliary ports, yielding a 5-port network whose scattering matrix can be determined analytically, see Eq. (19).
Using basic circuit analysis, as detailed in Appendix A, we readily obtain the $5\times 5$ scattering matrix of the circuit from Fig. 2B:
$\mathbf{S}^{\mathrm{SLC,T}}=\begin{bmatrix}0.25&0.25&-0.75&-0.25&0.5\\ 0.25&0.25&0.25&0.75&0.5\\ -0.75&0.25&0.25&-0.25&0.5\\ -0.25&0.75&-0.25&0.25&-0.5\\ 0.5&0.5&0.5&-0.5&0\end{bmatrix}.$ (18)
We have verified this result with three sanity checks, as detailed in Appendix A.
@@
$\mathbf{S}^{\mathrm{SLC,\pi}}=\begin{bmatrix}-0.25&0.25&-0.5&0.75&0.25\\ 0.25&-0.25&0.5&0.25&0.75\\ -0.5&0.5&0&-0.5&0.5\\ 0.75&0.25&-0.5&-0.25&0.25\\ 0.25&0.75&0.5&0.25&-0.25\end{bmatrix}.$ (19)
@@
The same exercise can be repeated for more complex ideal reconfigurable impedance networks such as the one from Fig. 2(b) in Ref. [7] but manual execution of the required circuit analysis rapidly becomes tiresome; already for the example from Fig. 2(b) in Ref. [7] the dimensions of the sought-after $\mathbf{S}^{\mathrm{SLC}}$ would be $14\times 14$. Therefore, for studies of large and/or realistic (as opposed to ideal) reconfigurable impedance networks, the numerical or experimental approaches described in the next two subsections are more suitable.